OpenStudy (anonymous):

Jethro is planting flowers in his garden. He wants the ratio of tulips to roses to be 4 to 2. Jethro wants to plant a total of 60 flowers. How many tulips should he plant?

OpenStudy (anonymous):

\(\cal\Huge\color{blue}{Welcome~To~OpenStudy!}\)

OpenStudy (anonymous):

The equivalent fractions to 4 to 2 are 2/1 = 4/2 = 6/3 = 8/4 = 10/5 = 12/6 = 14/7 = 16/8 = 18/9 = 20/10 = 22/11 = 24/12 = 26/13 = 28/14 = 30/15 = 32/16 = 34/17 = 36/18 = 38/19 = 40/20

OpenStudy (anonymous):

I was about to ask the same question, but I don't have to. Thank you @I_Need_Help_With_Home for showing us the equivalent fractions.